UT 2021 Oct 31 - Split Night PETS/PESI - D/E



Night Info:

  • PEPSI/PETS until UT 09:10 (02:10 local time)
  • Reconfigure 02:10 for LUCI-Bino (LUCI-2 Authorized)
  • Moon Rise: 01:47
  • Sunset: 17:35
  • Twilight Ends: 18:53
  • Twilight Starts: 05:13
  • Sunrise: 06:32

Night Plan:

  • LUCI-2 AGW-Transform - 1-1.5hr
    • thin or variable clouds, seeing < 1.2"
    • Transform Collect
    • Config: N30FieldStop, H-band filter, 4sec exposure time, N375 Camera
      • Potential lists:
        • Stone_C1.list - RA ~04hr Dec -01 - sets twilight
        • Stone_D_new_Oct27.list - RA ~04:30hr Dec -00 - thru twilight
        • Stone E (-90,-135,90, 0 PAs) - RA 07hr Dec +00 - 01:30hr-twilight
  • Active TMS testing - any conditions - 2-3hr

  • IF Conditions are too poor for LUCI-2 AGW Transform
    • Active TMS Testing
    • Arp 187 - Bino LUCI - 0.75hr (good conditions)
      • Script: /home/lbto/Service_Observing/2021B_D_E/LUCI/PoorWeather/LocalMergers
        • OT file: COBandhead_Mergers_V21.xml
        • sub-folder is Arp187 G200zJHK - special case HK 2.0 zJ 1.17 - in the CO Near-IR folder
        • finder chart (correctly oriented): Arp187_finder.jpeg
      • midnight until twilight
      • up to 2" seeing, 1-2 mag extinction (if cloudier double up number of exposures)
        • Fraternal is 1800sec zJ L1, HK L2
        • Mono is 3600sec HK, then 3600sec zJ
        • repeat script for poorest conditions
      • G200 zJ and HK 0.5" LS
      • Fraternal Setup (L1 zJ 1.17, L2 HK 2.0)
        • if bino-LUCI not available zJ and HK with mon-LUCI - HKspec highest priority
    • UGC 4079
      • same OT file as Arp 187, same conditions requirements.
      • Setup: G210 K-band CWL 2.3, 300sec dithers +/-45" along slit.
      • LUCI-1 and LUCI-2 versions included (if mono, repeat the script)
      • Finder chart is in directory
      • Target is located in the CO Near-IR folder
      • Arp 187 is higher priority than UGC 4079.
  • if LBC/TMS can't be used (or in addition to)
    • LUCI-Bino Photometric Standards (clear any seeing)
      • OT file: /home/lbto/Service_Observing/2021B_D_E/LUCI/LUCI_2021_Restart_V1.63.xml
      • Select the Photometric Standards Folder and pick ones visible
    • Arp 187 & UGC 4079 (see info above)

Night Log:

  • Prep for transform data collection
    • Use either Stone_D_new_27oct.list or a new list Stone_E_m135_m125_0.list (in ~/Service_Observing/2021B_E_D/LUCI). Stone D is setting at 2MST
    • In L2, put N30_FS and N3.75 camera
      • 0067 dark 2.5s x 1 x 1 LIR NORMAL
      • 0068 dark 4s x 1 x 1 LIR NORMAL

09:15 UT - Reconfiguring to bino-LUCI but only authorize LUCI2

09:25 UT - Josh is going to Stone_D and is pointed, now sending collimation. The collimation star was not found...

09:29 UT - Decided to switch to Stone_E since it is rising.

09:33 UT - Corrected pointing and sending collimation preset.

09:36 UT - Test 4-sec exposure through H. Oops - not yet on star.

09:38 UT - Send a monocular active preset to Stone_E TS=0, GS=0. Josh is homing AGW2.
  • Take test 4-sec exposure (0071)
  • 3-sec (0072)
  • 2-sec (2.5sec) 0073 is best.

09:46 UT - transform_collect started

index_# GS comments
0074 0  

09:50 UT - mask may not be centered ... it is very close (1019, 1021), but the star image being near the upper right (1050, 1060) on my test image threw me off. We had moved the guide probe and were not guiding, but little time elapsed.

09:54 - starting again - in fact the first image (0074) the star is closer to the center.

index GS Comments

0074 0 on-axis
0075 0 background
0076 70 really faint and star 0 near bottom left
0077 9 4 guide stars in the acq image, GCS chose the brightest, star 0 near bottom right of FS mask
0078 65 star 0 is near the bottom - wrong guide star chosen? there were two on the acq image
| 0079 | 21 | star 0 pretty well centered in mask
0080 85 stopping, GS not found --- star 0 is wobbling all over the place.

Is Stone_E too crowded? or is our pointing off?

10:05 Josh is going to check the pointing.

10:17 Checking count levels: 0081 H+4, 0082 H+2sec, 0083 FeII+4sec and 0084 is FeII+2sec

10:27 BrG and 2-sec. 0087 we've lost collimation. Try 0088 test image with 3-sec and BrG. But K is not the wavelength used for the DAR.

10:33 Try FeII+3 sec, star 0 has R=11.63. The star is just saturated and the spots are just visible - better when a 3-sec dark is subtracted but they are faint. I can see them with my eyes; hopefully the SW will find them.

The log for the run below is: /home/lbto/data/20211031/rfbg/transform_luci2.log

index GS comments
0088 0 star not saturated, spots visible
0089 - BG
0090 70  
0091 9 4 pix in core appear saturated, guider FWHM=1.2
0092 65 ok
0093 21 ok
0094 27 ok
0095 85 ok
0096 61 ok
0097 45 ok
0098 30 ok
0099 3 ok
0100 36 ok
0101 82 ok, but seeing blew up FWHM on image is 1.3" seeing may be trending up a bit but still 1.2 on average, 1.25-1.3 were the high values I saw on GCS.
0102 80 ok, but seeing not as good FWHM 1.05"
0103 77 FWHM 0.9 but elongated
0105 53 ok
0106 92 ok
0107 4 ok
0108 48 ok
0109 84 ok
0110 43 ok
0111 20 Seeing went back to subarcsec on the guider somewhere in the last few
0112 64 ok
0113 89 ok
0114 96 ok
0115 31 ok
0116 3 start of PA=-125 set, guide probe shadowing the upper right of the mask - it is very close
0117 84 ok, but star wobbled wrt 116
0118 21 ok, star wobble again
0119 20  
0120 89 out of range
0120 36 start=36 to skip 89, but I should have started at 32
0121 0 back on-axis
0122 31 start = 32
0123 77  
0124 70 ok
0125 82  
0126 36  
0127 0 on-axis
0128 3-sec dark  

04:55 We'll try Arp 187, however the conditions are not good - seeing is subarcsecond, but there is cirrus.

04:58 First, preset failed because I forgot to ask that LUCI1 be authorized, then it failed with a copointing error.

05:06 Checking and correcting pointing. First there was no star on the SX side. That was found, but now we are hunting for the star on DX. We are almost at 30 deg elevation and the object will be below 30 deg at 12 deg twilight.

05:13 Josh is going to an object higher up, back again to Stone_E, to check/correct pointing. After that, we can try the list to which I added a few stars (~/Service_Observing/2021B_D_E/LUCI/ Stone_Em135_m125_0_rev1.list, starting at 38.

05:25 We're really having to search for the star on DX again. LUCI-ARG is authorized on DX but LUCI on SX. It might have been that way since LUCI1 was authorized and that could explain the pointing problems.

Trying again with LUCI-SX and LUCI-DX but still not finding the star on DX. Josh is entering the last good values for bino LUCI and that worked.

06:00 pointed - but it is past 12-deg and probably too late to collimate and try to get a few more stars on Stone_E.

06:00 Josh is closing up.

-- BarryRothberg - 30 Oct 2021
Topic revision: r6 - 31 Oct 2021, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback